OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is odd and one die shows a 5? A 6 X 6 table of dice outcomes will help you to answer this question. Write fractions using the slash ( / ) key. Reduce fractions to their lowest terms

OpenStudy (slaaibak):

Tip: if the sum is odd, and the one number must be odd (a 5) the other must be even

OpenStudy (anonymous):

if one die shows 5, so another die will be the even so that they can have a odd sum. probability of one die shows 5 is 1/6 and probability of another die shows even is 3/6=1/2 so the final probability is (1/6)*(1/2)*2=1/6

OpenStudy (anonymous):

oh okay i get it know.thanks :)

OpenStudy (anonymous):

are u sure know that why i will multiply 2 at final probability?

OpenStudy (anonymous):

yes.thats how it shows on my apevx

OpenStudy (anonymous):

alright great

OpenStudy (anonymous):

or you can list them (5,2) (5,4) (5,6) (2,5) (4,5) (6,4) and \(\frac{6}{36}=\frac{1}{6}\)